please help me to find area of rectangle length 17ft 8in by width 11ft8in
Answer by TimothyLamb(4379)

a = (17*12 + 8) * (11*12 + 8)
a = (204 + 8) * (132 + 8)
a = (212) * (140)
a = 29680 sq.in
a = 29680 sq.in * (1 sq.ft / (12*12) sq.in ) = 206.11 sq.ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php